OpenStudy (anonymous):

These are part of our exam, i just want to share some questions with you..actually we are done with our exams..Find the coordinates of the point which divides the line segment from (-2,1) and (2,-3) in the ratio 3 to 4.

OpenStudy (anonymous):

Internal or external division?

OpenStudy (anonymous):

it is stated in the problem

OpenStudy (anonymous):

I don't think it's stated.

OpenStudy (anonymous):

ok i will give you a hint, from point(-2,1)

OpenStudy (anonymous):

Give me hint? Clarify the problem first.

OpenStudy (anonymous):

i think it would be an internal division wherein the point from (-2,1) would be three parts of the whole segment form.

OpenStudy (anonymous):

Okay so just use the formula ...

OpenStudy (anonymous):

of course, but you need to understand the problem first of what is asked?

OpenStudy (dumbcow):

The length between the 2 points is 4sqrt(2) The point (x,y) will be (3/7)(4sqrt2) from (-2,1) and (4/7)(4sqrt2) from (2,-3) using the distance formula \[(x+2)^{2} + (y-1)^{2} = \frac{288}{49}\] \[(x-2)^{2} + (y+3)^{2} = \frac{512}{49}\] of which a solution is \[(-.2857, -.714)\]